# Spacetime Gaussian Feature Splatting for Real-Time Dynamic View Synthesis

### 28 Dec 2023 v1

@@ Abstract
Novel view synthesis of dynamic scenes has been an intriguing yet challenging problem. Despite recent advancements, simultaneously achieving high-resolution photorealistic results, real-time rendering, and compact storage remains a formidable task. To address these challenges, we propose Spacetime Gaussian Feature Splatting as a novel dynamic scene representation, composed of three pivotal components. First, we formulate expressive Spacetime Gaussians by enhancing 3D Gaussians with temporal opacity and parametric motion/rotation. This enables Spacetime Gaussians to capture static, dynamic, as well as transient content within a scene. Second, we introduce splatted feature rendering, which replaces spherical harmonics with neural features. These features facilitate the modeling of view- and time-dependent appearance while maintaining small size. Third, we leverage the guidance of training error and coarse depth to sample new Gaussians in areas that are challenging to converge with existing pipelines. Experiments on several established real-world datasets demonstrate that our method achieves state-of-the-art rendering quality and speed, while retaining compact storage. At 8K resolution, our lite-version model can render at 60 FPS on an Nvidia RTX 4090 GPU.
@@ 1 Introduction
  • Extensive experiments on various real world datasets demonstrate that our method achieves state-of-the-art rendering quality and speed while keeping small model size. Our lite-version model enables 8K rendering at 60 FPS.
@@ 2 Related Work
Early approaches leverage image-based rendering techniques with proxy geometry/depth to sample novel views from source images [19, 29, 43, 32, 11, 95, 41, 13]. Chaurasia et al. [14] estimate a depth map to blend pixels from source views and employ superpixels to compensate for missing depth data. Hedman et al. [30] utilizes RGBD sensors to improve rendering quality and speed. Penner and Zhang [67] leverage volumetric voxels for continuity in synthesized views and robustness to depth uncertainty. Hedman et al. [31] learn the blending scheme with neural networks. Flynn et al. [24] combines multi-plane images with learned gradient descent.
@@
In recent years, neural scene representations have achieved great progress in novel view synthesis. These methods allocate neural features to structures such as volume [73, 57], texture [78, 16], and point cloud [1]. The seminal work of NeRF [61] proposes to leverage differentiable volume rendering. It does not require proxy geometry and instead uses MLPs to implicitly encode density and radiance in 3D space. Later on, numerous works emerge to boost the quality and efficiency of differentiable volume rendering. One group of methods focuses on improving the sampling strategy to reduce the number of point queries [63, 68, 4] or applies light field-based formulation [49, 74, 23, 82, 3, 76, 50]. Another group trades space for speed by incorporating explicit and localized neural representations [55, 93, 70, 77, 25, 88, 15, 62, 8, 34]. Among them, to improve model compactness, Instant-NGP [62] uses hash grid while TensoRF [15] utilizes tensor decomposition.
@@
A widely adopted setting for dynamic free-viewpoint rendering is using multi-view videos as input. Classic methods in this area include [37, 38, 90, 96, 18, 47, 46]. More recently, Broxton et al. [10] uses multi-sphere image as bootstrap and then convert it to layered meshes. Bansal et al. [5] separate static and dynamic contents and manipulate video with deep network in screen space. Bemana et al. [9] learn a neural network to implicitly map view, time or light coordinates to 2D images. Attal et al. [2] use multi-sphere representations to handle the depth and occlusions in 360-degree videos. Lin et al. [53] propose 3D mask volume to addresses the temporal inconsistency of disocclusions. Neural Volumes et al. [57] use an encoder-decoder network to encode images into a 3D volume and decode it with volume rendering. Lombardi et al. [58] enhance Neural Volumes with mixture of volumetric primitives. Extending static NeRF-related representations to dynamic scenes are also being actively explored [45, 83, 44, 75, 66, 84, 72, 26, 12, 4, 81, 36, 52]. DyNeRF [45] combines NeRF with time-conditioned latent codes to compactly represent dynamic scenes. StreamRF [44] accelerates the training of dynamic scenes by modeling the differences of consecutive frames. NeRFPlayer [75] decomposes scene into static, new and deforming fields and proposes streaming of feature channels. MixVoxels [81] represents scene with a mixture of static and dynamic voxels to accelerate rendering. HyperReel [4] utilizes sampling prediction network to reduce sampling points and leverages keyframe-based representation. K-Planes [26], HexPlane [12] and Tensor4D [72] factorize 4D spacetime domain into 2D feature planes for compact model size.
@@
Another line of research tackles dynamic view synthesis from monocular videos [48, 69, 27, 64, 79, 80, 86, 20, 65, 21, 28, 56, 51]. Under this setting, a single camera moves around in the dynamic scene, providing only one observed viewpoint at each timestep. To address the sparsity of supervision, priors on motion, scene flow or depth are usually introduced. In this work, we focus on the dynamic representation itself and only consider the setting with multi-view input videos.
@@
Recently, there are several work on this topic that are concurrent to ours [89, 59, 92, 91, 85, 87, 42, 35, 54]. 4K4D [89] combines 4D point clouds with K-Planes [26] and discrete image-based rendering, and uses differentiable depth peeling to train the model. Luiten et al. [59] models 4D scene with a set of moving 3D Gaussians, whose positions and rotations are discretely defined at each step. Their method demonstrates appealing results for 3D tracking, but its rendering quality is less favorable due to flickering artifacts. Yang et al. [92] leverages 4D Gaussians and 4D spherindrical harmonics for dynamic modelling. 4D Gaussians essentially represent motion with linear model. Comparatively, our polynomial motion model is more expressive, resulting in higher rendering quality. Our method also has higher rendering speed than their work. Yang et al. [91] and Wu et al. [85] prioritize on monocular dynamic view synthesis and employ deformation fields to deform a set of canonical 3D Gaussians. For multi-view videos setting, the performance of [91] is not extensively evaluated while [85] depicts inferior rendering quality and speed than ours.
@@ 4.1 Spacetime Gaussians
$\alpha_{i}(t)=\sigma_{i}(t)\exp\left(-\frac{1}{2}(\mathbf{x}-{\mu_{i}(t)})^{T}\Sigma_{i}(t)^{-1}(\mathbf{x}-\mu_{i}(t)\right),$ (6)
@@
We use a temporal radial basis function to represent the temporal opacity of an STG at any time $t$. Inspired by [39, 17] that use radial basis functions for approximating spatial signals, we utilize 1D Gaussian for the temporal opacity $P_{i}(t)$:
@@ 6.1 Neural 3D Video Dataset
The Neural 3D Video Dataset [45] contains six indoor multi-view video sequences captured by 18 to 21 cameras at $2704\times 2028$ resolution. Following common practice, training and evaluation are conducted in half resolution, and the first camera is held out for evaluation [45]. The number of frames is 300 for each scene.
@@ 6.2 Google Immersive Dataset
As shown in Tab. 2, our method outperforms NeRFPlayer and HyperReel in both speed and quality. Compared to HyperReel, our method is over 10 times faster in rendering speed. Although our PSNR is only 0.4 dB higher, the improvements in DSSIM and LPIPS are significant. When compared to NeRFPlayer, the margin is larger for all metrics. Visual comparisons are demonstrated in Fig. 4. Our method depicts sharper details and fewer artifacts than the others.
@@ 6.3 Technicolor Dataset
Tab. 3 shows the comparisons, where our method achieves noticeable gain in quality and speed. Please refer to supplementary material for visual comparisons on this dataset.
@@ 6.4 Ablation Study
Table 4: Ablation study of proposed components. Conducted on the Technicolor Dataset.
@@
To evaluate the effectiveness of proposed components, we conduct an ablation study in Tab. 4 using the Technicolor Dataset. Below we describe the configuration and performance of each ablation baseline.
Temporal Opacity.
“w/o Temporal Opacity” fixes the center and scale of the temporal radial basis functions in STGs during training. This variant suffers from a significant performance drop, revealing the importance of temporal opacity.
Polynomial Motion and Rotation.
“w/o Polynomial Motion” and “w/o Polynomial Rotation” fix the spatial position and rotation of STGs respectively. Both lead to a performance drop. Comparatively, motion is more important than rotation, which motivates us to use a lower-degree polynomial for rotation.
Feature Splatting.
“w/o Feature Splatting” uses the base RGB color $\mathbf{F}^{base}$ as the final color. It can be seen that there is a moderate drop in quality due to reduced ability to model view- and time-dependent appearance.
Guided Sampling of Gaussians.
“w/o Guided Sampling of Gaussians” does not encounter much performance drop in this dataset. The reason is that the scenes contain rich textures and can be well covered by SfM points. However, for other challenging scenes, guided sampling plays an important role (see Fig. 2 and supplementary material for examples).
Number of Frames used for Initialization.
We further analyzed the number of frames used for initialization in Tab. 5. Using fewer frames slightly downgrade quality, but also significantly reduces model size. It reveals that the compactness of our method can be further enhanced with a good selection of frames for initialization.
6.5 Limitation
@@ Appendix A Overview
  • Qualitative examples for the ablation on guided sampling in Appendix C.
  • Per-scene quantitative comparisons and more visual comparisons with other methods on the Neural 3D Video Dataset [45], Google Immersive Dataset [10] and Technicolor Dataset [71] in Appendix D.
@@
  • Real-time demos and dynamic comparisons in our video.
@@ Appendix B Comparisons with Concurrent Work
Appendix C Ablation on Guided Sampling
@@
We visualize the effects of guided sampling in Fig. 6. It can be seen that when without guided sampling, distant areas that are not well covered by SfM points will have very blurry rendering in both training and novel views. It reveals that it is challenging to pull Gaussians to these areas with gradient-based optimization and density control. On the other hand, with guided sampling applied, the rendering at these areas become much sharper for both training and novel views.
Appendix D More Results

### 04 Apr 2024 v2

@@ Abstract
Novel view synthesis of dynamic scenes has been an intriguing yet challenging problem. Despite recent advancements, simultaneously achieving high-resolution photorealistic results, real-time rendering, and compact storage remains a formidable task. To address these challenges, we propose Spacetime Gaussian Feature Splatting as a novel dynamic scene representation, composed of three pivotal components. First, we formulate expressive Spacetime Gaussians by enhancing 3D Gaussians with temporal opacity and parametric motion/rotation. This enables Spacetime Gaussians to capture static, dynamic, as well as transient content within a scene. Second, we introduce splatted feature rendering, which replaces spherical harmonics with neural features. These features facilitate the modeling of view- and time-dependent appearance while maintaining small size. Third, we leverage the guidance of training error and coarse depth to sample new Gaussians in areas that are challenging to converge with existing pipelines. Experiments on several established real-world datasets demonstrate that our method achieves state-of-the-art rendering quality and speed, while retaining compact storage. At 8K resolution, our lite-version model can render at 60 FPS on an Nvidia RTX 4090 GPU. Our code is available at https://github.com/oppo-us-research/SpacetimeGaussians.
@@ 1 Introduction
  • Extensive experiments on various real-world datasets demonstrate that our method achieves state-of-the-art rendering quality and speed while keeping small model size. Our lite-version model enables 8K rendering at 60 FPS.
@@ 2 Related Work
Early approaches leverage image-based rendering techniques with proxy geometry/depth to sample novel views from source images [20, 31, 46, 34, 11, 103, 44, 13]. Chaurasia et al. [14] estimate a depth map to blend pixels from source views and employ superpixels to compensate for missing depth data. Hedman et al. [32] utilize RGBD sensors to improve rendering quality and speed. Penner and Zhang [72] leverage volumetric voxels for continuity in synthesized views and robustness to depth uncertainty. Hedman et al. [33] learn the blending scheme with neural networks. Flynn et al. [26] combine multi-plane images with learned gradient descent. Wiles et al. [92] splat latent features from point cloud for novel view synthesis.
@@
In recent years, neural scene representations have achieved great progress in novel view synthesis. These methods allocate neural features to structures such as volume [78, 62], texture [84, 16], or point cloud [1]. The seminal work of NeRF [66] proposes to leverage differentiable volume rendering. It does not require proxy geometry and instead uses MLPs to implicitly encode density and radiance in 3D space. Later on, numerous works emerge to boost the quality and efficiency of differentiable volume rendering. One group of methods focuses on improving the sampling strategy to reduce the number of point queries [68, 73, 4] or applies light field-based formulation [53, 79, 25, 89, 3, 81, 54]. Another group trades space for speed by incorporating explicit and localized neural representations [60, 17, 83, 101, 75, 82, 27, 96, 15, 67, 8, 36]. Among them, to improve model compactness, Instant-NGP [67] uses hash grid while TensoRF [15] utilizes tensor decomposition.
@@
A widely adopted setting for dynamic free-viewpoint rendering is using multi-view videos as input. Classic methods in this area include [39, 40, 98, 104, 19, 49, 50, 51, 21]. More recently, Broxton et al. [10] use multi-sphere image as bootstrap and then convert it to layered meshes. Bansal et al. [5] separate static and dynamic contents and manipulate video with deep network in screen space. Bemana et al. [9] learn a neural network to implicitly map view, time or light coordinates to 2D images. Attal et al. [2] use multi-sphere representations to handle the depth and occlusions in 360-degree videos. Lin et al. [57, 58] propose 3D mask volume to addresses the temporal inconsistency of disocclusions. Neural Volumes [62] uses an encoder-decoder network to encode images into a 3D volume and decode it with volume rendering. Lombardi et al. [63] enhance Neural Volumes by decoding a mixture of dynamic geometric primitives from latent code and skipping samples in empty space for efficient ray marching. Extending static NeRF-related representations to dynamic scenes are also being actively explored [48, 90, 47, 80, 71, 91, 77, 28, 12, 4, 87, 38, 56, 88, 42]. DyNeRF [48] combines NeRF with time-conditioned latent codes to compactly represent dynamic scenes. StreamRF [47] accelerates the training of dynamic scenes by modeling the differences of consecutive frames. NeRFPlayer [80] decomposes scene into static, new and deforming fields and proposes streaming of feature channels. MixVoxels [87] represents scene with a mixture of static and dynamic voxels to accelerate rendering. HyperReel [4] utilizes sampling prediction network to reduce sampling points and leverages keyframe-based representation. K-Planes [28], HexPlane [12] and Tensor4D [77] factorize 4D spacetime domain into 2D feature planes for compact model size.
@@
Another line of research tackles dynamic view synthesis from monocular videos [52, 74, 29, 69, 85, 86, 94, 22, 70, 23, 30, 61, 55]. Under this setting, a single camera moves around in the dynamic scene, providing only one observed viewpoint at each timestep. To address the sparsity of supervision, priors on motion, scene flow or depth are usually introduced. In this work, we focus on the dynamic representation itself and only consider the setting of multi-view input videos.
@@
Recently, there are several work on this topic that are concurrent to ours [97, 64, 100, 99, 93, 95, 45, 37, 59]. 4K4D [97] combines 4D point clouds with K-Planes [28] and discrete image-based rendering, and uses differentiable depth peeling to train the model. Luiten et al. [64] models 4D scene with a set of moving 3D Gaussians, whose positions and rotations are discretely defined at each step. Their method demonstrates appealing results for 3D tracking, but its rendering quality is less favorable due to flickering artifacts. Yang et al. [100] leverage 4D Gaussians and 4D spherindrical harmonics for dynamic modelling. 4D Gaussians essentially represent motion with linear model. Comparatively, our polynomial motion model is more expressive, resulting in higher rendering quality. Our method also has higher rendering speed than their work. Yang et al. [99] and Wu et al. [93] prioritize on monocular dynamic view synthesis and employ deformation fields to deform a set of canonical 3D Gaussians. For multi-view videos setting, the performance of [99] is not extensively evaluated while [93] depicts inferior rendering quality and speed than ours.
@@ 4.1 Spacetime Gaussians
$\alpha_{i}(t)=\sigma_{i}(t)\exp\left(-\frac{1}{2}(\mathbf{x}-{\mu_{i}(t)})^{T}\Sigma_{i}(t)^{-1}(\mathbf{x}-\mu_{i}(t))\right),$ (6)
@@
We use a temporal radial basis function to represent the temporal opacity of an STG at any time $t$. Inspired by [41, 18] that use radial basis functions for approximating spatial signals, we utilize 1D Gaussian for the temporal opacity $\sigma_{i}(t)$:
@@ 6.1 Neural 3D Video Dataset
The Neural 3D Video Dataset [48] contains six indoor multi-view video sequences captured by 18 to 21 cameras at $2704\times 2028$ resolution. Following common practice, training and evaluation are conducted at half resolution, and the first camera is held out for evaluation [48]. The number of frames is 300 for each scene.
@@ 6.2 Google Immersive Dataset
As shown in Tab. 2, our method outperforms NeRFPlayer and HyperReel in both speed and quality. Compared to HyperReel, our method is over 10 times faster in rendering speed. Although our PSNR is only 0.4 dB higher, the improvements in DSSIM and LPIPS are significant. When compared to NeRFPlayer, the margin is larger for all metrics. Visual comparisons are shown in Fig. 4. Our method depicts sharper details and fewer artifacts than the others.
@@ 6.3 Technicolor Dataset
Tab. 3 shows the comparisons, where our method achieves noticeable gain in quality and speed. Please refer to supplementary material for visual comparisons.
@@ 6.4 Ablation Study
Table 4: Ablation study of proposed components. Conducted on all the five scenes from the Technicolor Dataset.
@@
To evaluate the effectiveness of proposed components, we conduct an ablation study in Tab. 4 using all the five scenes from Technicolor Dataset. Below we describe the configuration and performance of each ablation baseline.
Temporal Opacity. “w/o Temporal Opacity” fixes the center and scale of the temporal radial basis functions during training. This variant suffers from a significant performance drop, revealing the importance of temporal opacity.
Polynomial Motion and Rotation. “w/o Polynomial Motion” and “w/o Polynomial Rotation” fix the spatial position and rotation of STGs respectively. Both lead to a performance drop. Comparatively, motion is more important than rotation, which motivates us to use a lower-degree polynomial for rotation.
Feature Splatting. “w/o Feature Splatting” uses the base RGB color $\mathbf{F}^{base}$ as the final color. It can be seen that there is a moderate drop in quality due to reduced ability to model view- and time-dependent appearance.
Guided Sampling of Gaussians. “w/o Guided Sampling of Gaussians” does not encounter much performance drop in this dataset. The reason is that the scenes contain rich textures and can be well covered by SfM points. However, for other challenging scenes, guided sampling plays an important role (see Fig. 2 and supplementary material for examples).
Number of Frames used for Initialization. We further analyzed the number of frames used for initialization in Tab. 5. Using fewer frames slightly downgrade quality, but also significantly reduces model size. It reveals that the compactness of our method can be further enhanced with a good selection of frames for initialization.
6.5 Limitations
@@ Appendix A Overview
  • More ablation study in Appendix C.
  • Additional discussions in Appendix D.
@@
  • Per-scene quantitative comparisons and more visual comparisons with other methods on the Neural 3D Video Dataset [48], Google Immersive Dataset [10] and Technicolor Dataset [76] in Appendix F.
  • Real-time demos and dynamic comparisons in our video. Please refer to our website.
@@ Appendix B Comparisons with Concurrent Work
Appendix C More Ablation Study
C.1 Guided Sampling and Strategies of Adding Gaussians
@@
We visualize the effects of guided sampling in Fig. 6. It can be seen that when without guided sampling, distant areas that are not well covered by SfM points will have very blurry rendering in both training and novel views. It reveals that it is challenging to pull Gaussians to these areas with gradient-based optimization and density control. On the other hand, with guided sampling applied, the renderings at these areas become much sharper for both training and novel views. Note that the color tone difference in the bottom two rows is caused by inconsistent white balance in the training views of the scene, which makes each model have slightly different color tone in the novel view.
We also compare our guided sampling with two other strategies. The first one randomly adds Gaussians in the whole space and the second one adds a sphere of Gaussians near the far points of our guided sampling. As shown in Tab. 7 rows 2-5, our method has over $0.7$dB PSNR improvement.
C.2 Analysis on More Scenes
Tab. 7 rows 4-9 extend the ablation study in Tab. 4 of the main paper to additional scenes from the Neural 3D Video Dataset and the Google Immersive Dataset. We can see that our proposed components remain effective under various camera setup and scene content.
C.3 Polynomial Orders and Replacing Polynomials with MLP
In this experiment, we alter the polynomial orders $n_{p},n_{q}$ and replace the polynomials with MLPs.  Tab. 8 shows the results. Our choice of $n_{p},n_{q}$ and polynomials balances quality and storage.
Table 7: Ablation of guided sampling and other components. Conducted on the first 50 frames of Flame Salmon and 09_Exhibit scenes.
[…]
Add random Gaussians during init 27.72 0.0455 0.0787
Add a sphere of Gaussians during init 29.13 0.0381 0.0690
w/o Guided Sampling 27.48 0.0453 0.0921
Ours-Full 29.88 0.0373 0.0665
w/o Temporal Opacity 28.82 0.0376 0.0673
w/o Polynomial Motion 28.35 0.0406 0.0688
w/o Polynomial Rotation 28.69 0.0455 0.0690
w/o Feature Splatting 28.05 0.0448 0.0754
Table 8: Ablation of temporal function, polynomial orders, and features. Conducted on the first 50 frames of Theater and Sear Steak scenes.
[…]
Ours-Temporal-MLP 41.6 30.93 0.0428 0.0967
$n_{p}=1$ 33.4 32.24 0.0388 0.0823
$n_{p}=2$ 37.3 32.43 0.0379 0.0820
$n_{p}=4$ 44.3 32.61 0.0374 0.0809
$n_{q}=2$ 45.0 32.60 0.0377 0.0813
Ours-Full ($n_{p}=3,n_{q}=1$) 40.7 32.56 0.0376 0.0816
w/o $\mathbf{f}^{base}$ 31.9 31.83 0.0408 0.0959
w/o $\mathbf{f}^{dir}$ 38.5 32.03 0.0384 0.0849
w/o $\mathbf{f}^{time}$ 39.1 32.03 0.0392 0.0818
Random init $\mathbf{f}^{base}$ 36.3 32.14 0.0385 0.0841
Random init $\mathbf{f}^{dir}$ 40.7 32.22 0.0379 0.0827
Random init $\mathbf{f}^{time}$ 40.8 32.45 0.0378 0.0814
C.4 Feature Components
In this experiment, we ablate different features ($\mathbf{f}^{base}$, $\mathbf{f}^{dir}$ and $\mathbf{f}^{time}$) used in our full model. As shown in Tab. 8 rows 7-10 and Fig. 7, each component boosts rendering quality.
Figure 7: Ablation on feature components. Using all features produces the best visual quality.
Figure 8: Qualitative comparison of 3DGS [41], Ours-SH and Ours-Full. Conducted on the Flame Steak scene from the Neural 3D Video Dataset [48]. 3DGS is trained per-frame. Ours-SH denotes replacing our feature rendering with spherical harmonics rendering in 3DGS [41].
C.5 Initialization of Features
In our model, $\mathbf{f}^{base}$ and $\mathbf{f}^{dir}$ are initialized with the color of SfM points. $\mathbf{f}^{time}$ is initialized as zeros. The last three rows of Tab. 8 show an ablation of feature initialization, where our choice (Ours-Full) works better than random initialization.
C.6 Feature Rendering vs. Spherical Harmonics
In this experiment, we replace our full model’s feature rendering with spherical harmonics rendering, and refer to this baseline as Ours-SH.  Tab. 9 and Fig. 8 show that Ours-Full has better quantitative and visual quality while having smaller model size than Ours-SH. For fair comparisons of FPS, all methods use PyTorch implementation for rendering.
Table 9: Comparison with per-frame 3DGS and replacing the MLP in Ours-Full with SH. Conducted on the first 50 frames of Flame Salmon and Flame Steak scenes. Ours-SH uses SH of order 0 to 3, as in 3DGS.
[…]
3DGS 5100 135 700 29.76 0.0311 0.0486
Ours-SH 118 132 37 31.37 0.0276 0.0470
Ours-Full 37 145 35 31.66 0.0274 0.0467
C.7 Comparison with Per-Frame 3DGS
To validate the improvements of our method, we further compare with per-frame trained 3DGS. As shown in Tab. 9, our method has much smaller size and better rendering quality. Fig. 8 shows visual comparison.
C.8 Longer Video Sequence
In our experiments, following prior arts [80, 4], we train each model with 50-frame video sequence and arrange these models in series to render full-length sequences (typically 300 frames). In practice, this scheme can work for long videos at the cost of redundancy among models (e.g., static parts of the scene are repeatedly modeled). Our method also supports using a single model to represent more frames. Here, we conduct an experiment that directly trains our model with 300 frames on the Flame Salmon scene from the Neural 3D Video Dataset [48]. As shown in Tab. 10, compared to six 50-frame models in series, our single 300-frame model can reduce the per-frame training time and model size by around $80\%$ and $30\%$ respectively. At the same time, the rendering quality is comparable. This is attributed to our temporal opacity formulation so that complex long-sequence motion can be represented by multiple simpler motion segments.
Table 10: Performance of longer sequence per model on the Flame Salmon scene from the Neural 3D Video Dataset. We increase the training frames per model from 50 to 300. Longer sequence per model has smaller model size and shorter per-frame training time.
[…]
50 frames 6 12K 29.48 0.038 0.0224 0.063 20 1 300
300 frames 1 10K 29.17 0.037 0.0222 0.068 3.7 0.7 216
Appendix D Discussions
Our method is able to model shadows and ambient occlusions, as demonstrated in Fig. 7 and Fig. 8. For complex motion, our temporal opacity allows using multiple Gaussians where each one only needs to fit a shorter and less complex motion segment. Generally, the size of temporal RBF is small for fast-changing volumetric objects (e.g., flames) and large for static solid objects. Learned motion tends to be small for static objects and large for moving objects. Fig. 9 visualizes the temporal RBF and motion for an example scene. Note that our method does not apply additional regularization on motion.
For guided sampling, although it can alleviate the blurring in areas that are insufficiently covered by sparse point cloud, it cannot fully eliminate such artifacts. This is reflected in some challenging scenarios such as the far content outside windows in the Coffee Martini scene and the flame in the 02_Flames scene. The reason is that we do not have accurate depth of these areas, hence need to spawn new Gaussians across a depth range. However, these Gaussians may not cover the exact correct locations, and the ones near the correct locations may also be pruned during subsequent training. A possible solution would be to leverage the depth priors from learning-based depth estimation methods.
Figure 9: Visualization of trajectory across 25 frames. The background image is the ground truth at time $T=25$. The color of a trajectory denotes timestamp, where dark blue corresponds to $T=1$ and dark red corresponds to $T=25$. To visualize temporal opacity along a trajectory, we set the alpha channel value of a segment based on temporal opacity (excluding the spatial opacity term $\sigma_{i}^{s}$). We only show moving objects in the scene.
Appendix E Experiment Details
E.1 Baselines
Since the open source code of MixVoxels [87] does not contain the training config for MixVoxels-X, we use MixVoxels-L in our comparisons. We train HyperReel with their official code to generate visual examples. Note that the training time of HyperReel for each scene on the Technicolor Dataset is about 3.5 hours, while that of our full model on the Technicolor Dataset is only about 1 hour. For Dynamic 3DGS, its performance on the Neural 3D Video Dataset is not reported in their original paper. When applying their open source code to Neural 3D Video Dataset with default hyperparameters, the rendering quality is subpar. So we tune its hyperparameters to improve the performance on this dataset.
E.2 Camera Models
We use the original centered-undistorted camera model from 3DGS [41] for the Neural 3D Video Dataset. We implement the uncentered-undistorted camera model for the Technicolor Dataset. For the Google Immersive Dataset [10], to evaluate on the same distorted videos as [2, 80], we further adapt our method to fit the uncentered-distorted camera model with a differentiable image space warping, which maps perspective view to fish-eye distorted view. For the real-time demo, we retrain our models on the undistorted videos for simplicity. As there are black pixels in the undistorted images, we opt to use a mask to mask out the black pixels. Since image warping and masking are differentiable, our models can still be trained end-to-end.
E.3 Initialization
Following 3DGS [41], we use the sparse point cloud from COLMAP for initialization. Since the datasets provide camera intrinsics and extrinsics, we input them to COLMAP and call point triangulator to generate sparse points. The running time for point triangulator is much less than that of dense reconstruction. For the Theater, Train and Birthday scenes in Tab. 3 of the main paper and in Tab. 13, we only use 25 percent of SfM points from each frame (except the first frame whose SfM points are all used). The selection of SfM points is based on the distance between each point and its nearest neighbor. After sorting the distances, we keep the points with the longest distance to reduce redundancy. In the ablation study on the number of frames whose SfM points are used (Tab. 5 in the main paper), we use all the points in each sampled frames.
Features $\mathbf{f}^{base}$ and $\mathbf{f}^{dir}$ are initialized with the color of SfM points. $\mathbf{f}^{time}$ is initialized as zeros.
E.4 Density Control
During training, we conduct 12 times of cloning/splitting and over 50 times of pruning on the Technicolor dataset. Sparse points from multiple timestamps contain richer but more redundant information than sparse points from a single timestamp (or static points). Thus, after densification and guided sampling steps, we gradually prune Gaussians with small spatial opacity to keep the most representative Gaussians.
E.5 Guided Sampling
We uniformly sample from $s\times d$ to $7.5\times d$ with small random noise. $d$ is the max depth in a training view. $s$ is set as $0.7$ for most scenes.
E.6 Others
We apply sigmoid function to get the final RGB color for Neural 3D dataset and clamp function for the other two datasets in our full model. We use a linear form of the time variable and do not apply positional encoding on it. We use Nvidia RTX 3090 when reporting our rendering speed in the comparisons with other methods, and use Nvidia RTX 4090 in our real-time demos.
Appendix F More Results